A boat running at its' top speed takes 3.0 hours to travel 12 km down a river. It then takes 5.0 hours to return to its' starting place. What is the velocity of the river?
Answer by Alan3354(69443)

A boat running at its' [sic] top speed takes 3.0 hours to travel 12 km down a river. It then takes 5.0 hours to return to its' [sic] starting place. What is the velocity of the river?
-------------------
downstream speed = 12/3 = 4 km/hr
upstream speed = 12/5 = 2.4 km/hr
--------------
The boat's speed is the average = (4 + 2.4)/2 = 3.2 km/hr
Current is the difference = 4 - 3.2 = 0.8 km/hr
===================
PS There's no its' in English. It's its.